Calculate the mass (g) of magnesium oxide produced when 500 g of magnesium reacts with oxygen in the air.


2Mg + O2 → 2MgO


1
Expert's answer
2022-03-08T00:05:03-0500

If 500g of Mg react ;

Then , moles= 500/24= 20.8 moles

Mole ratio= 2.2

If 20.8 moles of magnesium react then 20.8 moles of magnesium oxide are produced.

Mass of MgO= 20.8× 40= 832g